1 A set of scores measuring aggression is nor- mally distributed with a mean equal to 23 and a standard deviation equal to 2.5. Find the proportion: la) To the left of x 19.0 b) To the right of x 25.5 (c) Between the mean and x 19.0 d] To the left of x 25.5 lel To the right of x 19.0
Answer #1

a)

for normal distribution z score =(X-μ)/σ
here mean=       μ= 23
std deviation   =σ= 2.5000
probability = P(X<19) = P(Z<-1.6)= 0.0548

b)

probability = P(X>25.5) = P(Z>1)= 1-P(Z<1)= 1-0.8413= 0.1587

c)

probability = P(19<X<23) = P(-1.6<Z<0)= 0.5-0.0548= 0.4452

d)

probability = P(X<25.5) = P(Z<1)= 0.8413

e)

probability = P(X>19) = P(Z>-1.6)= 1-P(Z<-1.6)= 1-0.0548= 0.9452